Calculus II, Spring 2014

Guide to Quiz 03

1. Section 14.2: Understand and memorize the meaning of limit and continuity for functions of two or more variables. Review the methods to check existence/non-existence of limit and continuity/discontinuity. Play attention to the examples and the homeworks on ‘Changing to Polar Coordinates’.

2. Section 14.3: Understand the meaning of ‘differentiable’ for functions of two or more variables. Understand the its connection with ’tangent plane’ of the graph of z = f(x, y) and why differentiability implies continuity. Can you find a function that is contunous but not differentiable at a point?

3. Section 14.4: Practice the Chain rule and implicit differentiation. Also pay attention to differentiating integrals.

4. Go through all home work problems. Think about what each problem is for, and whether you really leaned the point.